OpenStudy (anonymous):

please help, 21/4- 7/9-2/3

OpenStudy (anonymous):

First, make the denominators equal.

OpenStudy (anonymous):

ok done that

OpenStudy (anonymous):

So what do you have now?

OpenStudy (anonymous):

well i have 36 for my common denominator

OpenStudy (anonymous):

and i know i have to multiply but where do i borrow from

OpenStudy (anonymous):

No borrowing. For example first fraction, you're multiplying denominator by 9 to get 26. This means you just multiply the numerator by 9 too so you don't change the value. This would give you.\[\frac{21}4\times\frac99=\frac{189}{36}\]Just do the same with the rest.

OpenStudy (anonymous):

Oops, I meant 36 not 26.

OpenStudy (anonymous):

wait i typed it wrong its not 21 its 2 1/4

OpenStudy (anonymous):

\[\large \frac {21}{4}- \frac{7}{9}- \frac {2}{3}\]

OpenStudy (anonymous):

no 2 and 1/4

OpenStudy (anonymous):

\[\large \frac {21}{4} \times \frac {9}{9}- \frac{7}{9} \times \frac {4}{4}- \frac {2}{3} \times \frac {12}{12}\]

OpenStudy (anonymous):

\[\large 2\frac14=\frac94\]

OpenStudy (anonymous):

So just do what @dpalnc wrote down, just with 9 instead of 21.

OpenStudy (anonymous):

ok thank yall so much

OpenStudy (anonymous):

No problem :)